Grade 12th passPhysical Chemistry

One gm gm alloy of alluminum and mg mg heated with excess of Dil HCl .then h2 liberated at 1.102 and pressure 1 at step.

Tushar Anand
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
The ionic reaction equation for the reaction is :

1.) Mg(s) + 2H⁺ (aq) —> Mg²⁺(aq) + H₂(g)

Mole ratio is 1 : 2

2.) 2Al(s) + 6H⁺(aq) —> 2Al³⁺(aq) + 3H₂(g)

Mole ratio is 1 : 3

1 mole of H₂ = 22.4 litres at stp

1.12 / 22.4 = 0.05 moles.

Moles of Mg = 0.05 / 2 = 0.025 moles

Mass is :

0.025 × 24 = 0.6g

Moles of Al = 0.05 / 3 = 0.01667 moles.

Mass is :

0.01667 × 27 = 0.45 g

Composition of the alloy :

0.45 + 0.6 = 1.05g